# AllSpark: a multimodal spatiotemporal general model

### 31 Dec 2023 v1

@@ Abstract
RGB, infrared, multispectral, and other modal data fundamentally represent different observational approaches to the same geographic object. Therefore, leveraging multimodal data is an inherent requirement for comprehending geographic objects. However, for a long time, due to the high heterogeneity in structure and semantics among various spatiotemporal modal data, the joint interpretation of multimodal spatiotemporal data has been an extremely challenging problem. The primary challenge resides in striking a trade-off between the cohesion and autonomy of diverse modalities, and this trade-off exhibits a progressively nonlinear nature as the number of modalities expands. Inspired by the human cognitive system and linguistic philosophy, where perceptual signals from the five senses converge into language, we introduce the Language as Reference Framework (LaRF), a fundamental principle for constructing a multimodal unified model, aiming to strike a trade-off between the cohesion and autonomy among different modalities. Building upon this, we propose a multimodal spatiotemporal general artificial intelligence model, called AllSpark. Our model integrates thirteen different modalities into a unified framework, including one-dimensional (text, code), two-dimensional (RGB, infrared, SAR, multispectral, hyperspectral, tables, graphs, trajectory, oblique photography), and three-dimensional (point clouds, videos) modalities. To achieve modal cohesion, AllSpark uniformly maps diverse modal features to the language modality. In addition, we design modality-specific prompts to guide multi-modal large language models in accurately perceiving multimodal data. To maintain modality autonomy, AllSpark introduces modality-specific encoders to extract the tokens of various spatiotemporal modalities. And modal bridge is employed to achieve dimensional projection from each modality to the language modality. Finally, observing a gap between the model’s interpretation and downstream tasks, we designed task heads to enhance the model’s generalization capability on specific downstream tasks. Experiments indicate that AllSpark, without expert knowledge of the most spatiotemporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and trajectory compared to state-of-the-art models. Moreover, AllSpark showcases excellent adaptability in modalities like MSI, HSI, PointCloud, Table, Code, and Graph. AllSpark demonstrates the potential and possibility of constructing general artificial intelligence with a large language model. This approach contributes to the shift in research paradigm in spatiotemporal intelligence, transitioning from a modality-specific and task-specific paradigm to a general paradigm.
keywords:
Multimodal, Language, Spatiotemporal, General Model
1 Introduction
@@
Benefiting from the increasingly diverse observational methods for spatiotemporal scenes, geographic objects can be described using various spatiotemporal modalities such as RGB, infrared, SAR, multispectral, graphs, and spatiotemporal trajectories. Each modality provides unique information about different aspects of the geographic object. Analogous to the human process of perceiving and understanding the world through multiple modalities such as vision, hearing, and touch, joint interpretation of multimodal data is an inherent requirement for intelligent models to achieve cognition of geographic objects.
@@
However, due to the inherent differences in the mechanisms of each modality, various modalities are often highly heterogeneous in both structure and semantics. For example, in terms of structure, a table is composed of rows and columns. Point clouds are represented by three-dimensional coordinates along with various feature values. Text is constituted by sequences of words. In terms of semantics, RGB imagery reflects the electromagnetic characteristics of visible light bands emitted and reflected by geographic objects, whereas SAR imagery reflects the electromagnetic characteristics of microwaves emitted actively by radar for scattering from geographic objects.
@@
For a long time, constrained by the highly heterogeneity across various modalities as mentioned above, researchers have often developed specific methods based on prior assumptions related to a particular modality or designed multimodal approaches based on a few low-heterogeneity modalities. For instance, in single-modal research, C. R. Qi et al. proposed PointNetqi2017pointnet for the point cloud modality, emphasizing the invariance of point cloud data ordering and the significance of global and local features. For the language modality, Vaswani et al. introduced Transformervaswani2017attention , focusing on the long-range dependencies within word sequences. For the graph modality, Kipf and Welling proposed GCNkipf2016semi , based on the adjacency relationships between nodes in a graph. In multimodal research, fusion of optical and SAR imagery has been widely explored in both traditional and deep learning remote sensingkulkarni2020pixel . Text-visual models have also experienced rapid development in recent yearsradford2021learning ; lu2019vilbert ; kim2021vilt ; alayrac2022flamingo ; li2022blip . The diverse prior assumption associated with each modality have resulted in significant gaps between methods designed for different modalities, making it challenging to perceive and understand different modalities using a unified model.
@@
We believe that the key challenge in addressing this issue lies in striking a trade-off between the cohesion and autonomy of diverse modalities. Cohesion refers to the presence of mutually correlated shared information among modalities, while autonomy signifies the existence of unique information specific to each modality relative to others. If we simply project data from multiple modalities into a common representation space for fusion, emphasizing the cohesion between modalities, it may lead to the loss of modality-specific information, thereby diluting the essence of multimodal collaboration. On the contrary, if we excessively stress the autonomy between modalities, it may hinder the establishment of connections among them, limiting the model’s ability to simultaneously perceive multiple modalities. Moreover, as the number of modalities increases, balancing the cohesion and autonomy becomes progressively more challenging in a nonlinear manner.
@@
We observe that in the process of comprehending the world, humans integrate information from multiple modalities such as hearing, touch, smell, and vision. The concepts formed through the parsing of these modalities are ultimately converged in language. It is through language that humans engage in associating, reasoning, and expressive behaviors. In other words, language precisely encodes human perception and understanding of the world, providing clear definitions and meanings to abstract concepts from each modality. Inspired by this, we propose Language as Reference Framework (LaRF) as a fundamental principle for constructing multimodal models. As illustrated in Figure 1, the abstract concepts derived from each modality should align with language, allowing for joint interpretation in the unified representation space of language.
Figure 1: Language as Reference Framework (LaRF)
@@
Building upon this, we propose a multimodal spatiotemporal general artificial intelligence model, AllSpark, integrating thirteen different modalities into a unified framework, including one-dimensional (text, code), two-dimensional (RGB, infrared, SAR, multispectral, hyperspectral, tables, graphs, trajectory, oblique photography), and three-dimensional (point clouds, videos) modalities.
@@
To achieve modal cohesion, AllSpark uniformly maps diverse modal features to the language modality. In addition, we design modality-specific prompts to guide multi-modal large language models in accurately perceiving multimodal data.
@@
To maintain the autonomy between modalities, AllSpark introduce specific modal encoders for each modality to extract independent tokens. Given the high heterogeneity among modality data and modality encoders, there exists significant dimensional gap between the tokens of each modality and the language modality. To address this issue, we introduce modality bridge, a mechanism from Lynxzeng2023matters to accomplish the dimensional mapping from each modality’s tokens to the language tokens.
@@
Finally, considering the existing gap between the interpretation by the multimodal large language model and the specific downstream tasks, we have designed task heads for each downstream task to enhance the model’s generalization capability. It is noteworthy that, given the powerful interpretation capabilities of the multimodal large language model, we adhere to a lightweight design principle in both the modal encoders and task heads
@@
Experiments demonstrate that AllSpark, without the expert knowledge in most modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and spatiotemporal trajectories compared to state-of-the-art models. Specifically, in the RGB modality, AllSpark’s accuracy is only 0.84 lower than the optimal model, and in the trajectory modality, the ADE metric differs by only 0.07 compared to the optimal model. Additionally, AllSpark exhibits excellent adaptability in various other modalities, including point clouds, multispectral, hyperspectral, tables, graphs, and code. Theoretically, our proposed model has the potential for seamless extension to arbitrary number of modalities.
In a word, our contributions can be summarized as follows:
  • 1.
@@
    To our best knowledge, we first propose a unified multimodal spatiotemporal general model, AllSpark, successfully integrating 13 spatiotemporal modalities into a single model.
  • 2.
@@
    Inspired by the human cognitive system and linguistic philosophy, we propose the Language as Reference Framework (LaRF), offering a novel solution to balance cohesion and autonomy among multiple modalities. This principle provides a fresh approach to the multimodal unification.
  • 3.
@@
    We provide new possibilities for the paradigm of designing multimodal models in the spatiotemporal domain. MSTAGI demonstrates the potential and capabilities of building general model with large language models, thereby contributing to the shift in research paradigm in spatiotemporal intelligence, transitioning from a modality-specific and task-specific paradigm to a general paradigm.
  • 4.
@@
    Experiments indicate that AllSpark, without expert knowledge of the majority of spatiotemporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and trajectory compared to state-of-the-art models. Moreover, AllSpark showcases excellent adaptability in modalities like MSI, HSI, PointCloud, Table, Code, and Graph. Theoretically, our proposed model has the potential for seamless extension to arbitrary number of modalities.
2 Related Work
@@
Leveraging multimodal data is an inherent requirement for achieving cognitive recognition of geospatial objects. An ideal multimodal model should possess the capability to integrate all modalities for joint interpretation. Hence, a crucial trend in the research of intelligent methods in the spatiotemporal domain is the continual increase in the number of modalities for joint interpretation. Initially, early researchers often constructed single-modal expert models based on prior assumptions about a specific modality, achieving remarkable success within each respective modality. In recent years, with a deeper understanding of single-modal interpretation methods, numerous researchers have attempted to integrate several low-heterogeneity modalities to build multimodal interpretation approaches. However, as the number of modalities increases, the challenge of balancing cohesion and autonomy among the modalities becomes increasingly difficult. In the following, we recall the development of intelligent methods in the spatiotemporal domain from the perspective of the continually increasing number of modalities, and finally, we present the principles and approach of our proposed model, AllSpark.
2.1 Single-Modal Model
@@
For one-dimensional modalities, we focus on Code and Text. Code and Text are fundamentally language modalities, but due to the excellent characteristics of Code, such as strict syntax, unambiguous nature, and its ability to interact with machines, we treat it as a separate modality. Z. Feng et al. have pre-trained a model, CodeBERTfeng2020codebert , which facilitates the mutual transfer of information between Code and natural language modalities. For the Text modality, landmark contributions include Transformervaswani2017attention ,BERTkenton2019bert ,GPT series radford2018improving ; radford2019language ; brown2020language ; openai2023gpt4 , which have inspired a subsequent series of works.
@@
We categorize RGB, multispectral, hyperspectral, infrared, SAR, Table, Graph, Trajectory, and Oblique Photography as two-dimensional modalities.
@@
Among the two-dimensional modalities, the standard three-channel RGB image is one of the most common. For the RGB modality, the ResNethe2016deep proposed by K. He et al., based on the importance of visual global and local information, and the ViTdosovitskiy2020image introduced by A. Dosovitskiy et al., leveraging a global attention mechanism, represent two landmark contributions.
@@
An increase in the number of channels in images leads to the multispectral and hyperspectral modalities. B. Huang et al. proposed the STDCNNhuang2018urban , leveraging the characteristic of a higher number of bands in multispectral images to simultaneously model the global spatial and spectral properties of multispectral images. In comparison to multispectral images, hyperspectral images have even more bands, often reaching hundreds. Based on this, X. Yang et al. introduced the R-3D-CNNyang2018hyperspectral to further enhance the extraction of spectral features.
@@
Beyond an increase in the number of bands, attention has been directed towards images captured in specific spectral bands. For instance, infrared images, widely applied in fields such as agriculture and oceanography, have gained significant attention. H. Chen et al. proposed a shallow convolutional neural network, integrating a decision tree, to meet the demand for real-time intelligent interpretation of a large volume of infrared images in agricultural water resource monitoringchen2020deep . In the case of SAR images, formed by active microwave radar, S. Chen et al. introduced AConvNetchen2016target , a widely used fully convolutional neural network for SAR image intelligent interpretation.
@@
For oblique photography, 3D reconstruction aims to utilize single or multiple two-dimensional images from different angles to reconstruct a three-dimensional model. Y. Yao et al. addressed this task by proposing MVSNetyao2018mvsnet . Utilizing techniques such as three-dimensional convolution, the model establishes correlations between images from different perspectives, thereby reconstructing the three-dimensional model of the target from the features extracted from two-dimensional images.
@@
Beyond the visual domain, Table is one of the commonly used modalities for recording and expressing information. TabNetarik2021tabnet , proposed by S. Ö. Arik et al., employs a sequence attention mechanism to achieve feature selection in the Table modality, thereby enabling interpretable and more efficient learning.
@@
The trajectory modality reflects the temporal changes in the spatial positions of objects. A. Gupta et al. proposed Social GANgupta2018social based on the characteristic of trajectory multiplicity. This model combines historical trajectory information with social context to predict multiple plausible future outcomes.
@@
For the Graph modality, Kipf and Welling introduced the classic Graph Convolutional Network (GCN)kipf2016semi , which is based on the adjacency relationships between nodes in the graph. P. Velicˇkovic et al. proposed GATvelivckovic2017graph , which incorporates attention mechanisms into the Graph modality.
@@
Finally, we turn our attention to the three-dimensional modalities: Video and Point Cloud.
@@
The Video modality is formed by stacking images over time. A model proposed by A. Karpathy et al. extends the connectivity of CNN models over time. Additionally, it utilizes multi-resolution techniques to accelerate the training process, effectively reducing the impact of spatiotemporal redundancy on model efficiency in the Video modalitykarpathy2014large .
@@
The Point Cloud modality captures information about the position, shape, color, texture, and other aspects of three-dimensional objects. C. R. Qi et al. introduced the classic PointNetqi2017pointnet for the Point Cloud modality, emphasizing the importance of invariance through point data permutation and the significance of global and local features. W. Wu et al. extended convolution operations to three-dimensional point clouds with the introduction of PointConvwu2019pointconv .
2.2 Multi-modal Model

### 25 Feb 2024 v2

# AllSpark[Uncaptioned image]: A Multimodal Spatio-Temporal General Intelligence Model with Thirteen Modalities

@@ Abstract
RGB, infrared, multispectral, and other spatio-temporal modal data fundamentally represent different observational approaches for the same geographic object. Therefore, leveraging multimodal data is an inherent requirement for comprehending geographic objects. However, due to the high heterogeneity in structure and semantics among various spatio-temporal modal data, the joint interpretation of multimodal spatio-temporal data has long been an extremely challenging problem. The primary challenge resides in striking a trade-off between the cohesion and autonomy of diverse modalities. This trade-off becomes progressively nonlinear as the number of modalities expands. Inspired by the human cognitive system and linguistic philosophy, where perceptual signals from the five senses converge into language, we introduce the Language as Reference Framework (LaRF), a fundamental principle for constructing a multimodal unified model. Building upon this, we propose a multimodal spatio-temporal general artificial intelligence model called AllSpark. Our model integrates thirteen different modalities into a unified framework, including one-dimensional (text, code), two-dimensional (RGB, infrared, SAR, multispectral, hyperspectral, tables, graphs, trajectory, oblique photography), and three-dimensional (point clouds, videos) modalities. To achieve modal cohesion, AllSpark uniformly maps diverse modal features to the language modality by a modal bridge and a multimodal large language models. To maintain modality autonomy, AllSpark introduces modality-specific encoders to extract the tokens of various spatio-temporal modalities. Finally, observing a gap between the model‘s interpretability and downstream tasks, we designed modality-specific prompts and task heads to enhance the model’s generalization capability on specific downstream tasks. Experiments indicate that AllSpark, without expert knowledge of most spatio-temporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and trajectory compared to state-of-the-art models. Moreover, AllSpark shows excellent adaptability in various modalities, such as MSI, HSI, PointCloud, table, code, and graph. AllSpark demonstrates the potential and possibility of constructing general artificial intelligence with a large language model. LaRF and AllSpark contribute to the shift in the research paradigm in spatio-temporal intelligence, transitioning from a modality-specific and task-specific paradigm to a general paradigm.
Index Terms:
Spatio-temporal Data, Multimodal Machine Learning, Large Language Model, General Intelligence Model.
I Introduction
Figure 1: We propose ”Language as Reference Framework (LaRF)” as the fundamental principle for constructing multimodal models. Under the guidance of this principle, we introduced a multimodal general intelligence model, AllSpark, which integrates thirteen spatio-temporal modalities.
@@ I Introduction
Benefiting from the increasingly diverse observational methods available for spatio-temporal scenes, geographic objects can be described using various spatio-temporal modalities, such as RGB, infrared, SAR, multispectral, graph, and spatio-temporal trajectory data. Each modality provides unique information about different aspects of the geographic object. Analogous to the human process of perceiving and understanding the world through multiple modalities, such as vision, hearing, and touch, joint interpretation of multimodal data is an inherent requirement for intelligent models to achieve cognition of geographic objects.
@@
However, due to the inherent differences in the mechanisms of each modality, various modalities are often highly heterogeneous in both structure and semantics. For example, in terms of structure, a table is composed of rows and columns. Point clouds are represented by three-dimensional coordinates along with various feature values. A text is composed of sequences of words. In terms of semantics, RGB imagery reflects the electromagnetic characteristics of visible light bands emitted and reflected by geographic objects, whereas SAR imagery reflects the electromagnetic characteristics of microwaves emitted actively by radar for scattering from geographic objects.
@@
For a long time, constrained by the high heterogeneity across various modalities mentioned above, researchers have often developed specific methods based on prior assumptions related to a particular modality or designed multimodal approaches based on a few low-heterogeneity modalities. For instance, in single-modal research, C. R. Qi et al. proposed PointNet[1] for the point cloud modality, emphasizing the invariance of point cloud data ordering and the significance of global and local features. For language modality, Vaswani et al. introduced the transformer[2], which focuses on the long-range dependencies within word sequences. For the graph modality, Kipf and Welling proposed the GCN[3] based on the adjacency relationships between nodes in a graph. In multimodal research, the fusion of optical and SAR imagery has been widely explored in both traditional and deep learning remote sensing[4]. Moreover, visual-language models have undergone rapid development in recent years[5, 6, 7, 8, 9]. The diverse prior assumptions associated with each modality have resulted in significant gaps between methods designed for different modalities, making it challenging to perceive and understand different modalities using a unified model.
@@
We believe that the key challenge in addressing this issue lies in striking a trade-off between the cohesion and autonomy of diverse modalities. Cohesion refers to the presence of mutually correlated shared information among modalities, while autonomy signifies the existence of unique information specific to each modality relative to others. If we simply project data from multiple modalities into a common representation space for fusion, emphasizing the cohesion between modalities, this approach may lead to the loss of modality-specific information, thereby diluting the essence of multimodal collaboration. In contrast, if we excessively stress the autonomy between modalities, it may hinder the establishment of connections among them, limiting the model’s ability to simultaneously perceive multiple modalities. Moreover, as the number of modalities increases, balancing cohesion and autonomy becomes progressively more challenging nonlinearly.
@@
We observe that in the process of comprehending the world, humans integrate information from multiple modalities, such as hearing, touch, smell, and vision. The concepts formed through the parsing of these modalities ultimately converge in language. Humans engage in associating, reasoning, and expressive behaviours through language. In other words, language precisely encodes human perception and understanding of the world, providing clear definitions and meanings to abstract concepts from each modality. Inspired by this, we propose the Language as Reference Framework (LaRF) as a fundamental principle for constructing multimodal models. As illustrated in Figure 1, the abstract concepts derived from each modality should align with language, allowing for joint interpretation in the unified representation space of language.
@@
Building upon this, we propose a multimodal spatio-temporal general artificial intelligence model, AllSpark, that integrates thirteen different modalities into a unified framework, including one-dimensional (text, code), two-dimensional (RGB, infrared, SAR, multispectral, hyperspectral, tables, graphs, trajectory, oblique photography), and three-dimensional (point clouds, videos) modalities.
@@
To achieve modal cohesion, AllSpark uniformly maps diverse modal features to the language modality. In addition, we designed modality-specific prompts to guide multimodal large language models in accurately perceiving multimodal data.
@@
To maintain the autonomy between modalities, AllSpark introduces specific modal encoders for each modality to extract independent tokens. Given the high heterogeneity among modality data and modality encoders, a significant dimensional gap exists between the tokens of each modality and the language modality. To address this issue, we introduce a modality bridge, a mechanism from Lynx[10], to accomplish dimensional mapping from each modality’s tokens to the language tokens.
@@
Finally, considering the existing gap between the interpretability of the multimodal large language model and the specific downstream tasks, we design task heads for each downstream task to enhance the model’s generalization capability. Given the powerful interpretability capabilities of the multimodal large language model, we adhere to a lightweight design principle in both the modal encoders and task heads.
@@
Experiments demonstrate that AllSpark achieves competitive accuracy without expert knowledge in most modalities and utilizing a unified structure in modalities such as RGB and spatio-temporal trajectories compared to state-of-the-art models. Specifically, in the RGB modality, the accuracy of AllSpark is only 0.84 lower than that of the SOTA model, and in the trajectory modality, the average displacement error (ADE) metric differs by only 0.07 compared to that of the SOTA model. Additionally, AllSpark exhibits excellent adaptability in various other modalities, including point clouds, multispectral, hyperspectral, tables, graphs, and code. Theoretically, our proposed model has the potential for seamless extension to an arbitrary number of modalities.
In other words, our contributions can be summarized as follows:
@@
  • We first propose a unified multimodal spatio-temporal general model, AllSpark, that successfully integrates thirteen spatio-temporal modalities into a single model.
@@
  • Inspired by the human cognitive system and linguistic philosophy, we propose the language as reference framework (LaRF), which offers a novel solution to balance cohesion and autonomy among multiple modalities. This principle provides a fresh approach to multimodal unification.
@@
  • We propose new approaches for designing multimodal models in the spatio-temporal domain. AllSpark demonstrates the potential and ability to construct a general model with large language models, thereby contributing to the shift in the research paradigm in spatio-temporal intelligence, transitioning from a modality-specific and task-specific paradigm to a general paradigm.
@@
  • Experiments indicate that AllSpark, without expert knowledge of the majority of spatio-temporal modalities and utilizing a unified structure, achieves competitive accuracy in modalities such as RGB and trajectory information compared to state-of-the-art models. Moreover, AllSpark shows excellent adaptability in various modalities, such as MSI, HSI, PointCloud, table, code, and graph. Theoretically, our proposed model has the potential for seamless extension to an arbitrary number of modalities.
II Related Work
@@
Leveraging multimodal data is an inherent requirement for achieving cognitive recognition of geospatial objects. An ideal multimodal model should possess the capability to integrate all the modalities for joint interpretation. Hence, a crucial trend in the research of intelligent methods in the spatio-temporal domain is the continual increase in the number of modalities available for joint interpretation. Initially, early researchers often constructed single-modal expert models based on prior assumptions about a specific modality, achieving remarkable success within each respective modality. In recent years, with a deeper understanding of single-modal interpretation methods, numerous researchers have attempted to integrate several low-heterogeneity modalities to construct multimodal interpretation approaches. However, as the number of modalities increases, the challenge of balancing cohesion and autonomy among the modalities becomes increasingly difficult. In the following, we recall the development of intelligent methods in the spatio-temporal domain from the perspective of the continually increasing number of modalities, and finally, we present the principles and approach of our proposed model, AllSpark.
II-A Single-Modal Model
@@
For one-dimensional modalities, we focus on code and text. Code and text are both fundamental language modalities, but due to the excellent characteristics of code, such as strict syntax, unambiguous nature, and ability to interact with machines, code is treated as separate modality. Z. Feng et al. pretrained a model, CodeBERT[11], which facilitates the mutual transfer of information between code and natural language modalities. For the text modality, landmark contributions include a transformer[2], BERT[12], and the GPT series [13, 14, 15, 16], which have inspired subsequent series of works.
@@
We categorize RGB, multispectral, hyperspectral, infrared, SAR, table, graph, trajectory, and oblique photography as two-dimensional (2D) modalities.
@@
Among the two-dimensional modalities, standard three-channel RGB images are among the most common. For the RGB modality, the ResNet[17] proposed by K. He et al., which is based on the importance of visual global and local information, and the visual transformer (ViT)[18] introduced by A. Dosovitskiy et al., which leverages a global attention mechanism, represent two landmark contributions.
@@
An increase in the number of channels in images leads to multispectral and hyperspectral modalities. B. Huang et al. proposed the STDCNN[19], leveraging the characteristic of a greater number of bands in multispectral images to simultaneously model the global spatial and spectral properties of multispectral images. In comparison to multispectral images, hyperspectral images have even more bands, often reaching hundreds. Based on this, X. Yang et al. introduced the R-3D-CNN[20] to further enhance the extraction of spectral features.
@@
Beyond an increase in the number of bands, attention has been given to images captured in specific spectral bands. For instance, infrared images, widely applied in fields such as agriculture and oceanography, have gained significant attention. H. Chen et al. proposed a shallow convolutional neural network that integrates a decision tree to meet the demand for real-time intelligent interpretation of a large volume of infrared images in agricultural water resource monitoring[21]. In the case of SAR images formed by active microwave radar, S. Chen et al. introduced AConvNet[22], a widely used fully convolutional neural network for intelligent SAR image interpretation.
@@
For oblique photography, 3D reconstruction aims to utilize single or multiple two-dimensional images from various viewpoints to reconstruct a three-dimensional model. Y. Yao et al. addressed this task by proposing MVSNet[23]. By utilizing techniques such as three-dimensional convolution, the model establishes correlations between images from different perspectives, thereby reconstructing a three-dimensional model of the target from the features extracted from two-dimensional images.
@@
Beyond the visual domain, the table is one of the commonly used modalities for recording and expressing information. TabNet[24], proposed by S. Ö. Arik et al., employs a sequence attention mechanism to achieve feature selection in the Table modality, thereby enabling interpretable and more efficient learning.
@@
The trajectory modality reflects the temporal changes in the spatial positions of objects. A. Gupta et al. proposed the social generative adversarial networks (GAN)[25] based on the characteristic of trajectory multiplicity. This model combines historical trajectory information with social context information to predict multiple plausible future outcomes.
@@
For the graph modality, Kipf and Welling introduced the classic graph convolutional network (GCN)[3], which is based on the adjacency relationships between nodes in the graph. P. Veličković et al. proposed the GAT[26], which incorporates attention mechanisms into the graph modality.
@@
Finally, we turn our attention to three-dimensional modalities: video and point clouds.
@@
The video modality is formed by stacking images over time. A model proposed by A. Karpathy et al. extends the connectivity of CNN models over time. Additionally, multiresolution techniques are utilized to accelerate the training process, effectively reducing the impact of spatio-temporal redundancy on model efficiency in the video modality[27].
@@
The point cloud modality captures information about the position, shape, colour, texture, and other aspects of three-dimensional objects. C. R. Qi et al. introduced the classic PointNet[1] for the point cloud modality, emphasizing the importance of invariance through point data permutation and the significance of global and local features. W. Wu et al. extended convolution operations to three-dimensional point clouds with the introduction of PointConv[28].
II-B Multimodal Model